OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 15 meters and legs that measure 9 meters and 12 meters. What is the surface area of this prism if the height of the prism is 11 meters?

jimthompson5910 (jim_thompson5910):

What is the area of the triangle?

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

Area of triangle = (base*height)/2

OpenStudy (anonymous):

54

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

Now multiply that by the height of the prism (11) to get the volume of the prism

OpenStudy (anonymous):

594

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thats not an answer option

jimthompson5910 (jim_thompson5910):

oh I misread and thought they wanted the volume...one sec

OpenStudy (anonymous):

450 m2 486 m2 504 m2 612 m2

jimthompson5910 (jim_thompson5910):

The prism looks like this 

jimthompson5910 (jim_thompson5910):

There's a front face and a back face which are both triangles with the same area

jimthompson5910 (jim_thompson5910):

So far, we know that the front face is 54 sq meters, so the back face is also 54 sq meters

jimthompson5910 (jim_thompson5910):

In total, the front and back face add to 54+54 = 108 sq meters

jimthompson5910 (jim_thompson5910):

Now we do each lateral face Face1 = (Hypotenuse)*(Prism Height) Face2 = (Base)*(Prism Height) Face3 = (height of triangle)*(Prism Height)

jimthompson5910 (jim_thompson5910):

Making sense?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

So what's the area of face1?

OpenStudy (anonymous):

165?

jimthompson5910 (jim_thompson5910):

good, what about face2?

OpenStudy (anonymous):

99

jimthompson5910 (jim_thompson5910):

and face3?

OpenStudy (anonymous):

132

jimthompson5910 (jim_thompson5910):

now add up all the faces

OpenStudy (anonymous):

396

jimthompson5910 (jim_thompson5910):

that's the total lateral area, now add that to the area of the triangular faces

OpenStudy (anonymous):

which is?

jimthompson5910 (jim_thompson5910):

remember each triangular face is 54 square meters

jimthompson5910 (jim_thompson5910):

so add on 54+54 = 108

OpenStudy (anonymous):

504

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

can you help on another question I'm about to post?

jimthompson5910 (jim_thompson5910):

sure